OpenStudy (anonymous):

The coach throws a baseball to a player with a initial speed of 20 m/s at an angle of 45º with the horizontal. At the moment the ball is thrown, the player is 50 m from the coach. At what speed must the player run to catch the ball at the same height at which it was released?

OpenStudy (anonymous):

The answer is 3.19 m/s But I don't know how to get there... Can someone show me how to get this answer? Thanks.

OpenStudy (anonymous):

find the time and distance traveled when its height is zero (back to the same height it started at). then find the velocity the player must have to cover the distance from his x position (x=50) to the x location where the ball 'lands' in the time it takes the ball to travel to that point.

OpenStudy (anonymous):

solve for t: \[height: 0 = 20*\sin 45 *t - \frac{ g t ^{2} }{ 2 }\] using t, solve for x \[distance: x = 20*\cos 45 *t \] distance - 50 is how far he has to run, and the time you found is how long he has to cover that distance

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Why is height = 0? We don't know height.

OpenStudy (anonymous):

"to catch the ball at the same height at which it was released" delta h =0

OpenStudy (anonymous):

Oh ok. Thanks.

OpenStudy (anonymous):

get the right answer now?